# Generative Bid Shading in Real-Time Bidding Advertising

### v1

@@ 1. Introduction
Online advertising is now a primary way for companies to attract users and build their brands. Real-Time Bidding (RTB) is the key system in programmatic advertising, allowing auctions in milliseconds. When a user visits a website, the Supply-Side Platform (SSP) sends an ad request to the Ad Exchange (ADX), where Demand-Side Platforms (DSPs) bid based on request information. The highest bidder wins the chance to show their ad. The auction way has changed a lot due to transparency issues in auctions and compatibility challenges with emerging technologies (AppNexus, 2017; Getintent, 2017; Sluis, 2017). Most major platforms, following Google (Bigler, 2019), have switched from Second-Price Auctions (SPA) to First-Price Auctions (FPA). In SPA, advertisers usually bid through their actual value, but in FPA or non-ideal SPA settings, this can cause them to pay much more than necessary. For example, if an advertiser bids \$5 while the second-highest bid is only \$2, they would overpay up to \$3. By dynamically adjusting bids toward the minimum winning price, bid shading (Crespi and Sexton, 2005) significantly optimizes cost structures and directly improves resource allocation efficiency in the trillion-dollar digital advertising market.
Figure 1. Overview of two-stage bid shading process.
Bid shading is a censored problem (Huang et al., 2024) because of the inaccessibility of the minimum winning price. Some methods assume an uncensored scenario in which the ADX will return the minimum winning price regardless of the auction outcome, allowing them to fit an optimal shading ratio using linear regression. However, this assumption is unrealistic in real-world DSP environments: in FPA, the minimum winning price is unobservable regardless of whether the bid wins or loses, resulting in no ground truth for bid shading. Most existing bid shading strategies adopt a two-stage framework, as illustrated in Fig. 1. The first stage is machine learning (ML), which estimates the bid landscape (i.e., the winning price or winning rate distribution) to compute the winning rate and expected cost for varying bids. The second stage is operations research (OR), which employs optimization algorithms (e.g., bisection search, golden section search) to identify the bid that maximizes surplus, using the ML predictions as input.
Figure 2. The winning rate and surplus curves derived from Meituan’s non-ideal SPA data show that surplus first increases and then decreases as bids increase, but do not strictly follow a standard unimodal pattern. As a result, two-stage methods are challenging in finding the global optimum. In contrast, GBS takes a broader view of the generative space, allowing it to escape the local optimum effectively.
Despite its widespread adoption, the two-stage method faces two significant limitations. First is the non-convex surplus profile: existing methods typically strongly assume that the surplus optimization curve is unimodal (Pan et al., 2020; Zhou et al., 2021), thereby enabling search algorithms to determine the optimal shading ratio. However, in real-world scenarios, winning rate curves are often non-smooth, and cost curves in non-ideal SPA exhibit significant fluctuations. This results in a non-convex surplus profile, causing optimization algorithms to converge to local optima, as illustrated in Fig. 2 (a). Second, cascading errors from the ML+OR pipeline: the sequential workflow introduces cascading errors. In the precise mathematical operations of the OR stage, errors originating from the ML stage can be further amplified. Consequently, even theoretically optimal OR algorithms may fail to achieve practical optimality. An ideal approach is to evaluate all bid candidates in real time to identify the globally optimal bid. However, such an exhaustive search is computationally infeasible in a millisecond-latency RTB system.
A recent study (Gong et al., 2023) has attempted to generate the shading ratio in an end-to-end manner. However, their use of negative log-surplus as the loss function in supervised learning encounters the same limitations as two-stage methods when faced with non-convex surplus curves: convergence to local optima. Furthermore, their simple discriminative models struggle to meet the high-precision requirements of bidding scenarios. Traditional generative algorithms, such as conditional variational autoencoders (CVAE (Sohn et al., 2015)), are constrained by predefined prior distributions (e.g., Gaussian), which limits their performance ceiling in shading ratio generation. To overcome distributional assumptions and mitigate outlier interference, existing direct scalar-output methods often discretize continuous values (Sun et al., 2024; Lin et al., 2023). However, this approach fails to leverage conditional dependencies between discrete intervals; discretization is only reflected in label definitions without mutual interdependence (e.g., 0.3 is closer to 0.4 than 0.2), and the lack of error-correction mechanisms leads to suboptimal results. Additionally, DSPs optimize holistic revenue during bidding, causing model performance to be heavily influenced by bid price distributions. This introduces significant data selection bias, as the training data fails to represent the real action space of bids.
@@
In this paper, we propose a novel end-to-end Generative Bid Shading (GBS) framework that leverages an autoregressive generative model and reward-preference alignment to address the challenges above from a global perspective, as illustrated in Fig. 2 (b). GBS comprises two main components:
1). An autoregressive generative model: Inspired by the success of large language models (LLMs) in recommendation and regression tasks, we adopt an encoder-decoder autoregressive architecture to generate shading ratios. We construct a data-driven vocabulary of primitive tokens, encoding any shading ratio scalar as a token sequence. The model decomposes shading ratio generation into a conditional and sequential modeling process, where each step predicts a ratio component, and the output at each time step serves as input for the next. This method effectively captures dependencies across multiple prediction steps and progressively approximates the shading ratio in a residual manner. The token sequence generation enables coverage of a broader value range and adaptation to complex distributions. To accelerate model convergence and ensure stability, we employ teacher forcing during pre-training and introduce gumbel softmax to guarantee gradient backpropagation and consistency between training and inference.
2). A reward preference alignment system: To overcome the limitations of traditional approaches, we introduce a preference alignment reward system and train the model using reinforcement learning in the generated space. We first propose a channel-aware hierarchical dynamic network (CHNet) as the reward model to capture fine-grained features for precise bid landscape modeling. Next, we introduce two preference alignment modules to guide learning. Specifically, surplus optimized alignment maximizes the final surplus by ensuring that bids for obtainable requests are not overbidding, while exploration utility alignment incorporates bid exploration into the bid shading system, alleviating sample selection bias and optimizing long-term surplus. Finally, we employ the group relative policy optimization (GRPO) for reinforcement learning to achieve efficient and stable policy updates.
In summary, our contributions are as follows:
  • We identify the suboptimality in two-stage bid shading methods and propose a generative framework that autoregressively generates shading ratios with error correction capabilities from a global perspective. To the best of our knowledge, this work represents the first implementation of generative bid shading in large-scale online systems.
  • We design a CHNet that captures fine-grained features for the shifting bidding landscape. Based on this, we introduce two reward preference alignment modules: one maximizes immediate surplus by avoiding overbidding; the other mitigates data selection bias via strategic bid exploration. This paradigm enhances generalization across diverse business scenarios and extensibility.
@@
  • Through extensive experiments on both offline and online A/B tests, we show that our GBS significantly outperforms the state-of-the-art methods. GBS has been deployed in Meituan DSP and serves billions of bid requests daily, demonstrating industrial viability and efficiency.
@@ 2. Related Work
Generative Model and Preference Alignment. With the rapid development of large language models (LLMs), generative models are increasingly used in areas such as generative recommendation (Deng et al., 2025) and search (Guo et al., 2025). Preference learning, including Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), aligns model outputs with human preferences but is inefficient. Direct Preference Optimization (DPO) (Rafailov et al., 2023) and its variants (Azar et al., 2024; Chowdhury et al., [n. d.]) address this by directly adjusting model policies using preference data. However, they face challenges like sample sparsity in recommendation and advertising. Group Relative Preference Optimization (GRPO) (Shao et al., 2024) uses explicit reward functions and group-based evaluation to enhance decision-making and training efficiency. In this paper, we present the first exploration of applying the generative model to bid shading, leveraging reward systems to guide the model beyond the constraints of traditional methods.
@@ 3.2. Generative Model Preprocessing
Specifically, we introduce a vocabulary $\mathcal{V}=\left\{w_{j}\right\}_{j=1}^{V}$ where each element $w_{i}$ represents a predefined scalar value, analogous to tokens in the language model. The vocabulary embedding matrix is denoted as ${E}\in\mathbb{R}^{V\times D}$, where $V$ consists of the vocabulary size, and $D$ is the dimension of the token embeddings. Each shading ratio can be encoded as a sequence of tokens $s_{i}=\left\{s_{i}^{1},\ldots,s_{i}^{L}\right\}$, with $L$ indicating the length of the sequence. This process is known as tokenization. In contrast to generative recommendation systems that index directly based on ID, we propose a label decoding function $g(\cdot)$ to determine the actual scalar value from the token ID output by the model, thereby deriving the shading factor $\alpha_{i}=g(s_{i})=\sum_{l=1}^{L}g(s_{i}^{l})\in\mathbb{R}$.
@@ 4. Method
In this section, we provide a detailed explanation of the proposed GBS. First, we introduce the end-to-end autoregressive generative model, which maps the shading factor through the output token sequence. Next, we present the reward preference alignment system, which employs a channel-aware hierarchical dynamic network as the reward model and utilizes GRPO to guide policy learning via two reward alignment modules.
@@ 2. Negative Log-Loss of expected surplus for winning data $N_{+}$:
(9) $L_{\text{surplus}}=\sum_{i=1}^{N_{+}}\mathbb{E}((v_{i}-C(v_{i}*\alpha_{i}))\cdot Wr(v_{i}*\alpha_{i}\mid x_{i})),$
@@ 4.2.1. Channel-Aware Hierarchical Dynamic Network (CHNet)
In real-world advertising systems, varying ADXs/SSPs employ opaque and diverse billing methods, leading to significant differences in the bid landscape across channels. Inspired by multi-scenario approaches (Gao et al., 2023), we propose a channel-aware hierarchical dynamic network (CHNet) for fine-grained channel information modeling, as shown in Fig. 4. CHNet utilizes a hierarchical structure with an explicit channel-oriented layer and multiple implicit channel-oriented layers, sequentially modeling coarse-grained explicit information and fine-grained implicit information. For processing feature matrices222for simplification, request $i$ is not separately denoted in this section, we adhere to the method outlined in Section 4.1, with separate processing for channel feature $e^{c}$ and others $e^{o}$.
@@ 5. Experiments
In this section, we conduct extensive experiments to verify the performance of GBS, where implementation details are detailed in Appendix C.
@@ 5.1.1. Datasets
We evaluate our method on both the public and real-world industrial datasets, with detailed statistics presented in Table 1. The public iPinYou (Liao et al., 2014) data are derived from the SPA setting, where the advertiser’s bid is used as the actual value in the bid request, and the paid price is regarded as the winning price. The dataset is randomly split into training and testing sets with a ratio of 7/3. Our private dataset is collected from one month of bid sampling data from Meituan DSP and is partitioned similarly. Notably, only the testing set contains minimum winning price, which is used for evaluation.
@@ 5.2. Overall Performance
CAVE 51.42% 96,981,112 36.45% 26,698,064
Post-CAVE 57.14% 107,781,128 36.63% 26,830,032
@@ 5.5. Performance on Online System (Meituan)
To evaluate the online performance of GBS, we deployed it on the Meituan DSP. We conducted a rigorous A/B test over two weeks, and the engineering architecture is shown in Fig. 7. We allocated 30% of the traffic to GBS, while the baseline method was an online two-stage bid shading approach, which utilizes the golden section search in the operations research stage. We model certain channels as a non-ideal SPA scenario, consistent between baseline and GBS. We use four key online metrics for evaluation: Return On Investment (ROI), Cost Per Mille (CPM), Cost Per Click (CPC), and Inference Time (IT).

### v2

@@ 1. Introduction
Online advertising has emerged as a cornerstone of the digital economy, with Real-Time Bidding (RTB) serving as the fundamental infrastructure for programmatic media buying. In this ecosystem, Demand-Side Platforms (DSPs) compete in auctions to purchase impressions on behalf of advertisers. Under the conventional RTB paradigm, these high-frequency auctions were typically implemented using a second-price rule, under which the highest bidder won the impression but paid a settlement price equal to the second-highest bid. This effective mechanism can incentivize truthful bidding by aligning bidders’ dominant strategy with submitting offers that reflect their underlying valuation of each impression, thereby mitigating the risk of systematic overpayment relative to prevailing competitive conditions.
However, the practical application of second-price auctions has been increasingly challenged by a lack of transparency regarding settlement prices and an inability to adapt to complex, emerging auction environments (Bigler, 2019; AppNexus, 2017; Getintent, 2017; Sluis, 2017). Recently, the industry has shifted to First-Price Auctions (FPA), where the winner pays their own bid, implying that bidding one’s full valuation generally reduces surplus and results in systematic overpayment. In an idealized setting where the market-clearing (minimum winning) price for a given bid request was observable, an advertiser could bid exactly that price to win the impression at the lowest possible cost. However, because this clearing price is latent and should be inferred under substantial uncertainty, advertisers instead rely on bid shading (Crespi and Sexton, 2005; Huang et al., 2024), which strategically discounts the submitted bid below the raw valuation. By calibrating this discount to trade off win probability against surplus preservation, bid shading can substantially reduce acquisition costs and improve budget efficiency, thereby enhancing resource allocation in the trillion-dollar digital advertising market.
Figure 1. Overview of bid shading process.
The prevailing approach to bid shading typically adopts a two-stage pipeline (Fig. 1): it first estimates the bid landscape, and then uses an operations-research (OR) module to maximize surplus via search algorithms (e.g., bisection search (Pan et al., 2020), golden section search (Zhou et al., 2021)). However, this decoupled design is inherently brittle. Its rigidity leads to a key limitation: the final bid can be suboptimal because the optimizer relies on simplifying assumptions and is sensitive to estimation errors. Many OR solvers assume that the surplus as a function of the bid is unimodal (Pan et al., 2020; Zhou et al., 2021), so that simple search procedures can reliably locate the optimum. In practice, the surplus curve in RTB is often non-convex and may contain multiple local peaks, causing search algorithms to converge to a local optimum rather than the global one. As illustrated in Fig. 2, the surplus curve initially increases and then decreases, but does not strictly adhere to a standard unimodal pattern. In principle, one could evaluate all candidate bids online to find the best bid, but exhaustive evaluation is computationally infeasible under millisecond-level latency constraints. Moreover, because the workflow is sequential, errors in the first-stage estimation are directly carried into the second-stage optimization and can be further magnified, preventing the system from achieving practical optimality within strict latency budgets.
To address these issues, we explore a generative paradigm that mitigates local optima by combining diverse candidate generation with relative evaluation. In this paradigm, we sample multiple bid candidates from a generative model and evaluate them jointly, which approximates a broader search than single-point optimization. We further compute relative advantages among candidates, so learning is driven by comparisons rather than fragile absolute targets, reducing the chance of being misled by local extrema. As illustrated in Fig. 2, conventional search methods frequently stagnate in local optima when facing non-convex objective functions, while GBS adopts a broader view of the generative space, allowing it to effectively converge to the global optimum. However, implementing this idea is non-trivial with existing generative architectures. Methods such as CVAE (Sohn et al., 2015) often rely on rigid predefined priors, while discretization-based approaches (Sun et al., 2024; Lin et al., 2023) fail to leverage conditional dependencies between discrete intervals, as they treat these intervals as isolated labels without mutual interdependence. Moreover, training data collected from historical winning bids is inherently biased and covers only a limited portion of the counterfactual action space, further restricting effective learning.
Figure 2. The winning rate and surplus curves derived from Meituan’s non-ideal second-price auctions data.
@@
Building upon the above insights, we incorporate autoregressive modeling (Ma et al., 2024) and reinforcement-based alignment (Deng et al., 2025) to propose a novel end-to-end framework, Generative Bid Shading (GBS). This approach leverages a global perspective to address the limitations of traditional two-stage methods, utilizing a generative architecture to optimize bid shading. GBS comprises two main components:
1) An autoregressive generative model: To address the limitations of rigid priors and the neglect of dependencies between discrete intervals, we propose an autoregressive model. By decomposing shading ratios into sequences of discrete tokens, our method treats generation as a conditional sequential process. This design enables the model to effectively capture high-order dependencies between token values, facilitating precise approximation through residual refinement. To support this, we employ teacher forcing and gumbel-softmax to ensure stable convergence and end-to-end differentiability.
2) A reward preference alignment system: We propose a reinforcement learning-based preference alignment system anchored by the Channel-aware Hierarchical Dynamic Network (CHNet), a high-precision reward model. Policy optimization is driven by two distinct alignment mechanisms: a surplus optimization alignment that maximizes immediate returns by curbing overbidding on obtainable requests, and an exploration utility alignment that integrates strategic bid exploration to counteract sample selection bias. Furthermore, we employ Group Relative Policy Optimization (GRPO) to guarantee efficient and stable policy updates throughout the training process.
Crucially, this design inherently circumvents the issues of local sub-optimality and cascading errors that typically plague two-stage approaches. In summary, our contributions are as follows:
  • We are the first to formally identify the inherent limitations of the conventional two-stage bid shading paradigm. We demonstrate that the decoupling of estimation and optimization leads to cascading errors, while the reliance on unimodal assumptions often results in local optimality within non-convex surplus landscapes, thereby fundamentally constraining the performance upper bound of existing methods.
  • We propose a novel end-to-end generative paradigm designed to transcend traditional limitations. This framework synergistically integrates a novel autoregressive generative model, a high-precision reward model (CHNet), and dual preference alignment mechanisms, which simultaneously maximize immediate surplus and mitigate data selection bias, achieving superior scalability and generalization across diverse business scenarios.
@@
  • Through extensive experiments on both offline and online A/B tests, we show that our GBS significantly outperforms the state-of-the-art methods. GBS has been deployed in Meituan DSP and serves billions of bid requests daily. To the best of our knowledge, this work represents the first implementation of generative bid shading in large-scale online systems.
@@ 2. Related Work
Generative Model and Preference Alignment. With the rapid development of large language models (LLMs), generative models are increasingly used in areas such as generative recommendation (Deng et al., 2025) and search (Guo et al., 2025). Preference learning, including Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), aligns model outputs with human preferences but is inefficient. Direct Preference Optimization (DPO) (Rafailov et al., 2023) and its variants (Azar et al., 2024; Chowdhury et al., ) address this by directly adjusting model policies using preference data. However, they face challenges like sample sparsity in recommendation and advertising. Group Relative Preference Optimization (GRPO) (Shao et al., 2024) uses explicit reward functions and group-based evaluation to enhance decision-making and training efficiency. In this paper, we present the first exploration of applying the generative model to bid shading, leveraging reward systems to guide the model beyond the constraints of traditional methods.
@@ 3.2. Generative Model Preprocessing
Specifically, we introduce a vocabulary $\mathcal{V}=\left\{w_{j}\right\}_{j=1}^{V}$ where each element $w_{i}$ represents a predefined scalar value, analogous to tokens in the language model. The vocabulary embedding matrix is denoted as ${E}\in\mathbb{R}^{V\times D}$, where $V$ consists of the vocabulary size, and $D$ is the dimension of the token embeddings. Each shading ratio can be encoded as a sequence of tokens $s_{i}=\left\{s_{i}^{1},\ldots,s_{i}^{L}\right\}$, with $L$ indicating the length of the sequence. This process is known as tokenization. In contrast to generative recommendation systems that index directly based on ID, we propose a label decoding function $g(\cdot)$ to determine the actual scalar value from the token ID output by the model, thereby deriving the shading ratio $\alpha_{i}=g(s_{i})=\sum_{l=1}^{L}g(s_{i}^{l})\in\mathbb{R}$.
@@ 4. Method
In this section, we provide a detailed explanation of the proposed GBS. First, we introduce the end-to-end autoregressive generative model, which maps the shading ratio through the output token sequence. Next, we present the reward preference alignment system, which employs a channel-aware hierarchical dynamic network as the reward model and utilizes GRPO to guide policy learning via two reward alignment modules.
@@ 2. Negative Log-Loss of expected surplus for winning data $N_{+}$:
(9) $L_{\text{surplus}}=-\sum_{i=1}^{N_{+}}\mathbb{E}\left(\log\left((v_{i}-C(v_{i}\cdot\alpha_{i}))\cdot Wr(v_{i}\cdot\alpha_{i}\mid x_{i})\right)\right),$
@@ 4.2.1. Channel-Aware Hierarchical Dynamic Network (CHNet)
In real-world advertising systems, varying media employ opaque and diverse billing methods, leading to significant differences in the bid landscape across channels. Inspired by multi-scenario approaches (Gao et al., 2023), we propose a channel-aware hierarchical dynamic network (CHNet) for fine-grained channel information modeling, as shown in Fig. 4. CHNet utilizes a hierarchical structure with an explicit channel-oriented layer and multiple implicit channel-oriented layers, sequentially modeling coarse-grained explicit information and fine-grained implicit information. For processing feature matrices222for simplification, request $i$ is not separately denoted in this section, we adhere to the method outlined in Section 4.1, with separate processing for channel feature $e^{c}$ and others $e^{o}$.
@@ 5. Experiments
In this section, we conduct extensive experiments to verify the performance of GBS.
@@ 5.1.1. Datasets
We evaluate our method on both the public and real-world industrial datasets, with detailed statistics presented in Table 1. The public iPinYou (Liao et al., 2014) data are derived from the second-price auctions setting, where the advertiser’s bid is used as the actual value in the bid request, and the paid price is regarded as the winning price. The dataset is randomly split into training and testing sets with a ratio of 7/3. Our private dataset is collected from one month of bid sampling data from Meituan DSP and is partitioned similarly. Notably, only the testing set contains minimum winning price, which is used for evaluation.
@@ 5.2. Overall Performance
CVAE 51.42% 96,981,112 36.45% 26,698,064
Post-CVAE 57.14% 107,781,128 36.63% 26,830,032
@@ 5.5. Performance on Online System (Meituan)
To evaluate the online performance of GBS, we deployed it on the Meituan DSP. We conducted a rigorous A/B test over two weeks, and the engineering architecture is shown in Fig. 7. We allocated 30% of the traffic to GBS, while the baseline method was an online two-stage bid shading approach, which utilizes the golden section search in the operations research stage. We model certain channels as a non-ideal second-price auctions scenario, consistent between baseline and GBS. We use four key online metrics for evaluation: Return On Investment (ROI), Cost Per Mille (CPM), Cost Per Click (CPC), and Inference Time (IT).